OpenStudy (swaf11):

6/5 divided by blank = 6 a.)1/5 b.)30 c.)1/3 d.)5

OpenStudy (f_jayyy):

A because the lower part of the fraction can go in 6 times

OpenStudy (f_jayyy):

so it is "A"

OpenStudy (swaf11):

oh, so like multiply the denominator by the numinator

OpenStudy (gabbyalicorn):

6/5/x=6 6/5/x=6 Multiply both sides by 5x. 6=30x 30x=6(Flip the equation) 30x/30=6/30(Divide both sides by 30) x=15 Check answers. (Plug them in to make sure they work.) x=15(Works in original equation) Answer x=1/5 A

jabez177 (jabez177):

Don't give out direct answers please, @F_Jayyy. :) We are here to guide the people to their answers. :)

OpenStudy (f_jayyy):

i didnt jabez look at my first comments

OpenStudy (anonymous):

I agree with Gabby :D

OpenStudy (anonymous):

Also "6/5 divided by blank = 6" could be represented as 6/5/x=6

jabez177 (jabez177):

Actually you did, @F_Jayyy, you said "A is the one because so and so." You are supposed to guide them, not tell them and explain it. :)